### v1

# Universality in the velocity jump in the crack propagation observed for food-wrapping films for daily use

Abstract
The velocity jump found in the crack propagation for rubbers has been a powerful tool for developing tough rubber materials. Although it is suggested by a theory that the jump could be observed widely for viscoelastic materials, the report on a clear jump is very limited and, even in such a case, reproducibility is low, except for elastomers. Here, we use a mundane food-wrapping film as a sample and observe the crack propagation velocity with pulling the sample at a constant speed in the direction perpendicular to the crack. As a result, we find the jump occurs at a critical strain with high reproducibility. Remarkably, the plot of the crack-propagation velocity as a function of strain can be collapsed onto a master curve by an appropriate rescaling, where the master curve is found to be universal for change in the pulling speed and in the sample height. The result reveals a key parameter for the jump is the strain, suggesting the existence of a small length that governs the deformation along the crack. The present study sets limitations on future theories and opens an avenue for the velocity jump to become a tool for developing a wide variety of tough polymer-based materials.
Introduction — To develop tough polymer-based materials are crucial in our modern life. Accordingly, the fracture of polymer based materials has been studied intensively [1, 2, 3, 4, 5, 6], but remains an active area of research. One of the principal topics in the study of polymer fracture is the crack propagation, which can be studied through various methods: the cyclic tensile test [7, 8], trouser test [9, 10], static tensile test [11, 12], and so forth [13, 14].
Among a myriad of polymer-based materials, rubber is one of ancient matters human beings have been exploited and is used on a daily base in various forms from automobile tires to anti-vibration system for buildings. In the study of fracture of rubber or elastomer, the velocity jump has been discussed systematically [11, 12] and more recent years extensively with an increasing expectation for toughening [15, 16, 17, 18, 19].
Conventionally, the jump is observed in experiments performed in a static experiment performed under a static boundary condition. In the static experiment, a film sample with the width much larger than the height is stretched in the height direction, and with a fixed stretch a cut is introduced at one of the free edges. The cut triggers crack propagation in the direction of width and the propagation speed reaches a constant value, which defines the crack-propagation velocity for the fixed strain. The velocity as a function of strain increases as the strain increases, which may be intuitively natural. Remarkably, the velocity jumps typically 10000 times at a critical strain. This dramatic phenomena is called the velocity jump.
In 2016, a theoretical explanation was proposed based on a simple viscoelastic model, which suggests the jump can be observed for a wide range of viscoelastic materials [18]. The jump was reproduced in numerical simulation [17], and the physical picture for the origin of the velocity jump proposed in Ref[18], the glass transition at the crack tip, is strengthened by a study performed under a collaboration among theoretical, numerical and experimental groups [20].
Since the suggestion that the jump can be observed for viscoelastic materials other than rubber, some attempts have been made to observe the jump in other polymer-based materials, such as gel [21, 22]. However, observation of a clear jump other than elastomers has been limited. One such example is known for a porous sheet made of polypropylene, but the reproducibility of the experiment was low [23]. Although a clear jump, of the order of four orders of magnitude, was observed for any samples used, the quantities characterizing the jump, such as the critical strain and the velocities just before and after the jump, were clearly dependent on samples.
As a matter of fact, the velocity jump was not observed for the porous polypropylene, when the crack propagation was observed under the static boundary condition [24]. However, the jump was observed in a dynamic experiment, in which the crack propagation was observed under a dynamic boundary condition [23]: we first introduce a short crack at one edge of the sample, and then start stretch the sample at a fixed speed to observe the crack propagation [25]. The reason the jump was not observed under the static experiment may be that the sample is relaxed during the preparation for giving a fixed stretch, while such effects of relaxation can be reduced in the dynamic experiment. This is one of the advantages of the dynamic experiment over the static one.
Another advantage of the dynamic experiment is efficiency in time and cost. In the static experiment, we need many samples to obtain a velocity curve as a function of strain, because we need to break one sheet to obtain a single point on the curve. On the contrary, in the dynamic experiment, we can obtain a full curve with a single sheet: we can measure the velocity as a function of strain during crack propagation because the strain changes as we stretch the sample.
Here, we use a mundane food-wrapping film as a sample and observe the crack propagation under the dynamic boundary condition. As a result, we find the jump occurs at a critical strain with high reproducibility. The crack-tip position as a function of time is obtained using two cameras, one for capturing the slow crack propagation before jump and the other with high frame rate for the fast crack propagation after jump. The plot of the crack-propagation velocity as a function of strain thus obtained shows a little sample dependence: for a fixed pulling velocity and a fixed sample size, we obtained quantitatively similar curves for different samples. Furthermore, we find that the velocity curve is independent from the pulling velocity but dependent on the sample height, while they can be collapsed onto a master curve by a rescaling: the master curve thus obtained is universal for change in the pulling speed and the height of the sample. The result reveals a key parameter for the jump is the strain, suggesting the existence of a small length that governs the deformation along the crack. The present study sets limitations on future theories and opens an avenue for the velocity jump to become a tool for developing a wide variety of tough polymer-based materials.
[…]
From the strain-stress curve obtained at the room temperature at the strain rate $\simeq 0.02$ to 0.07 s-1, the elastic modulus is around 500 MPa and increases with the strain rate. From rheological measurements, the sample exhibits a glass transition at the time scale around tens of milliseconds at the room temperature, where the elastic modulus at the glass phase $\simeq 5$ GPa drops to a smaller value. See Appendix for further details on mechanical and rheological properties.
Figure 1: (a) Experimental setup for the dynamic experiment. (b) Illustration for the long and thin tape-like area along the crack path.
Experimental — As illustrated in Fig. 1, we cut a sheet sample from a role of width $W$ (=150=150 mm) and set the sheet bottom to the fixed clamp and the sheet top to the mobile clamp so that the sample height at the zero strain becomes $L$ (=10=10 to 25 mm), i.e., the distance between the bottom and top clamps is set to $L$ at the zero strain. We introduce a short line crack of length $a$ at one of the free edge, before we start stretching the mobile clamp at a constant speed $U$ (=0.5=0.5 to 2 mm/s). The mobile clamp is connected to a slider (EZSM6D040K, Oriental Motor) controlled by a controller unit (ESMC-K2, Oriental Motor) via a wire. The strain rate $U/L$ is in the range 0.025 to 0.2 s-1 in the velocity jump experiment. All experiments are performed at the room temperature $\simeq 25$ C.
Some of the stress-strain curve given in Appendix are obtained by the same setup used for the velocity jump experiment. But, in this case, the mobile clamp is not connected to the slider system directly via wire but to a digital force gauge (FGP-50, NIDEC-SHIMPO) mounted on the slider system via a wire.
We record the crack propagation using two cameras. In order to capture the slow propagation before the jump, we record a video at 60 frames per second (fps) with a digital camera (D800E, Nikon) with a lens (AF-S Micro NIKKOR 60mm 1:2.8G ED, Nikon). In addition, to capture the fast propagation after the jump, we record images at 32000 fps with a high-speed camera (FASTCAM Mini UX100, Photron) with a lens (AF-S NIKKOR 20mm 1:1.8G ED, Nikon).
To clearly observe the crack propagation, before we set the film sample to the clamps, we paint the long and thin tape-like area along the crack path as shown in Fig. 1 (b), whose area is $Wl$ with $l$ typically 1 cm, with an oil-based white marker (Mackee Paint Marker, Zebra). But, later, to see the effect of painting, we performed experiments by painting the whole area (=WL=WL) with a different white water-based pigment marker (POSCA Extra Thick Square Core, UNI). In the following, we mainly discuss the data obtained from samples with the whole area painted and compare the data obtained from partially painted films. In the two different experiments, we also change the length of the initial crack $a$: in the partial painting experiment we set $a=10$ mm, but in the whole painting experiment, we set $a=L$. By comparing the two experiments, we can gain insight into the effect of both painting and the initial crack length $a$.
Figure 2: (a) Crack tip position $X$ vs. time $t$ before the jump at $X=X_{c}$ and $t=t_{c}$. (b) Crack tip position $\Delta X$ vs. time $\Delta T$ after the jump, where $\Delta X=X-X_{c}$ and $\Delta T=t-t_{c}$. The three sets of the data are obtained from different three samples for $U=1$ mm/s and $L=15$ mm.
Crack propagation — From the images obtained from the digital camera, we track the crack tip position $X$ as a function of time $t$, before the jump, as in Fig. 2 (a), where we showed the results from three samples obtained for the same set of $(U,L)$. The three sets of the results are well overlapped, demonstrating high reproducibility of the slow propagation. As for the fast propagation after the jump, we track the crack position $\Delta X(t)=X(t)-X(t_{c})$ as a function of time $\Delta T=t-t_{c}$ with $t=t_{c}$ corresponding the time of jump, as in Fig. 2 (b), where we show the results from three samples obtained for the same set of $(U,L)$. The three sets of the results are again well overlapped, demonstrating high reproducibility even for the fast propagation.
Figure 3: (a) Velocity $V$ vs. strain $\varepsilon$ for a fixed $L$, exhibiting a collapsed straight region on the log-log scale before the jump and a collapsed point just after the jump. (b) $V$ vs. $\varepsilon$ for a fixed $U$. (c) All the data in (a) and (b) on a rescaled horizontal axis $\varepsilon L$, showing $V$ is a function of $\varepsilon L$ in the straight region and at the points just before and after the jump. (d) $V$ vs. $\varepsilon L$ obtained for various parameters by a different way of painting to clearly observe the crack propagation: partial painting (see the text for the details).
Velocity jump — In Fig. 3 (a), we show the crack-propagation velocity as a function of strain for various pulling speed $U$ at a given sample height $L$ (=15=15 mm), which shows a clear jump. The velocity jump of three orders of magnitude occurs at a critical strain $\varepsilon_{c}\simeq 0.1$ with the velocity just before the jump $V_{B}\simeq 200$ mm/s and that just after the jump $V_{A}\simeq 300$ m/s. The curve shows a clear straight region in the log-log plot before the jump. The straight region and the velocities just before and after the jump, $V_{B}$ and $V_{A}$, are superimposed well even though the data with different $U$ are shown, which demonstrates with a high reproducibility the crack-propagation dynamics leading the jump is independent of the pulling speed $U$.
In Fig. 3 (b), we show the same relation for various sample height $L$ at a given pulling speed $U$ (=2=2 mm/s), which again shows a clear jump. We see that the velocities just before and after the jump, $V_{B}$ and $V_{A}$, are almost the same, while the dependence of the critical strain on $L$ is recognized.
In Fig. 3 (c), we show all the data in Fig. 3 (a) and (b) with rescaling the horizontal axis from $\varepsilon$ to $\varepsilon L$. The plot clearly demonstrates that all the data in the region from the straight region to the point after the jump collapse onto a master curve, which is universal for changes in $U$ and $L$. The slope $\alpha$ of the straight region is $\sim 4$.
To see the effect of painting and the initial crack length $a$, we show the data obtained from partially painted films in Fig. 3 (d), with a straight line on the log-log plot with a slope $\alpha\sim 6$. Although we see slight changes in $V_{B}$ and $V_{A}$ as well as in the slope of the straight region from Fig. 3 (c), the basic features are unchanged: (1) $V=F(\varepsilon L)$ except for the small strain region with $F(\varepsilon L)\simeq(\varepsilon L)^{\alpha}$, where the master curve $F(x)$ is universal even if we change $U$, $L$, and $a$ in limited ranges, although the slope $\alpha$ is affected probably by painting. (2) $V_{B}$ and $V_{A}$ are also almost independent from $U$, $L$, $a$, and the way of painting.
To obtain velocity of the slow crack propagation in the above analysis, we fit the relation between the crack tip position $X$ and time $t$ with the sixth-order polynomial of $t$ and take the derivative of the polynomial with respect to $t$. We checked that the results do not exhibit significant changes if we use the fifth-order polynomial except for the small strain region before the straight region on the log-log plot. The velocity of the fast crack propagation was obtained by a linear fitting.
Discussion — We find the crack propagating velocity is given as a function of $\Delta L=\varepsilon L$, i.e., the amount of the absolute stretch $\Delta L$, regardless of the sample height $L$. We interpret this unexpected result as follows. There is a characteristic microscopic length scale $\simeq l^{\ast}$ near the crack and practically only the region near the crack tip characterized by $l^{\ast}$ is stretched. In other words, the strain near the crack tip is not governed by $\Delta L/L$ but by $\varepsilon^{\ast}=\Delta L/l^{\ast}$ and the velocity is a function of this effective strain $\varepsilon^{\ast}$. One possible candidate for the scale $l^{\ast}$ may be defined by the characteristic time scale for the glass transition $t_{0}$ ($\simeq$ tens of milliseconds) we obtain from the rheology experiment and the velocity just before the jump $V_{B}\simeq 200$ mm/s: $l^{\ast}=V_{B}t_{0}\simeq$ a few mm. In this respect, the experiment with smaller $L$ with $L<l^{\ast}$ is interesting, although technically challenging.
The crack propagation velocity has been frequently plotted as a function of the fracture energy $G$ [27, 28, 29, 30], which scales as $\varepsilon^{2}L$ for liner elastic materials. This is the case in most of previous studies on the velocity jump (e.g., Ref[16]). However, we confirmed that, if we re-scale the horizontal axis not to $\varepsilon L$ but to $G\sim\varepsilon^{2}L$ in Fig. 3 (c) and (d), the collapse to a master curve is not observed anymore (they become scattered as in Fig. 3 (b)). Note that, in the present case, the critical strain at the jump is relatively small, up to which $G$ can be obtained with a reasonable precision under the assumption of a linear stress-strain curve (see Appendix).
The elastic modulus and the density of the present sample are $E\simeq 500$ MPa and $\rho\simeq 1000$ kg/m3 at the room temperature. The corresponding velocity of elastic wave is $v\simeq\sqrt{E/\rho}\simeq 500$ m/s, which gives a practical maximum of the crack propagation speed and is comparable with the velocity after the jump $V_{A}\simeq 300$ m/s: After the jump, the velocity reaches the maximum speed. This is in contrast with the case of elastomers, where $V_{A}$ is a few orders of magnitude slower than the velocity of the elastic wave $v\simeq 50$ m/s evaluated for the elastic modulus of the rubbery regime $E\simeq$ a few MPa (e.g., Ref[16]). Because of this, for elastomers we observe a ”fast-mode” in the velocity range from $V_{A}$ to $v$, in which $V$ scales with some power of $\varepsilon$. However, it could be possible that we observe velocity increase after the jump even in the present case if the sample width is much longer, or we perform the static experiment with minimizing the preparation time for relaxation of the set samples.
The analytical theory is based on a viscoelastic model characterized the following parameters [18]: (1) a rubber elastic modulus $E_{R}$ and a glassy one $E_{G}=\lambda E_{R}$ where $\lambda$ is typically $10^{3}$ for elastomers, (2) the sample height $L$ and a microscopic parameter $l_{0}=L/N$, (3) a viscosity $\eta$, and (4) a critical strain $\varepsilon_{c}$ for rupture. The theory predicts the normalized crack-velocity $V/V_{0}$ with $V_{0}\sim(l/\eta)\sqrt{E_{R}}$ as a function of the normalized energy density $w/w_{0}$ with $w\sim E_{R}\varepsilon^{2}$ and $w_{0}$ $\sim E_{R}\varepsilon_{c}^{2}/N$. Since $w/w_{0}\sim N\varepsilon^{2}/\varepsilon_{c}^{2}\sim\varepsilon^{2}L$ and the fracture energy $G\sim\varepsilon^{2}L$, the theory suggests that $V$ is given as a function of $G$ for a given material: $V=F(G)$. In addition, the theory predicts $V_{B}\sim\sqrt{\lambda}V_{0}$ and $V_{A}\sim\sqrt{N}V_{0}$ as for the characteristic velocities and predicts the energy density at the jump as $w_{j}\sim\lambda w_{0}$, which leads to a prediction for the strain at the jump: $\varepsilon_{j}^{2}\sim(E_{G}/E_{R})/N$.
However, a close comparison of the theory with the present experiment is difficult. This is because, while the theory is based on the single relaxation from the glass phase to the rubber phase, the present sample shows two relaxations: first for the glass phase and the second for the crystalline phase (see Appendix). Our experiments suggest, even in such a case, the glass transition, a common property in both cases of the theory and the present experiment, is the key for the velocity jump.
Conclusion — We found a clear velocity jump in a highly reproducible manner for a polymer film we use on a daily base under the dynamic experiment. The velocity - strain relation can be collapsed onto a master curve after rescaling, which is universal for the stretching speed and the sample size in limited ranges. Our study suggests that the velocity jump can be widely observed in various polymer films if they exhibit a glass transition in rheology measurements, which underscores the potential of the dynamics crack propagation experiment for developing tough polymers.
Acknowledgments — We thank Professor Takashi Uneyama (Nagoya University) for discussion on rheological properties of the sample. We are grateful to Anton Paar Japan K.K. for helpful technical support concerning rheological measurements. This work was supported by JSPS KAKENHI Grant Number JP19H01859 and JP24K00596.
Figure 4: Stress vs. Strain. (a) Dependence on the stretching direction. (b) Dependence on the sample size. (c) Dependence on the stretching speed. (d) Magnified version of (c).
Figure 5: Complex modulus and $\tan\delta$ as a function of temperature at $f=1$ and $10$ Hz.
Appendix: Mechanical Responses
Stress vs. Strain —
In Fig. 4, we show the stress-strain curve measured by two methods, one using a rheometer (MCR 302, Anton Paar) and the other using a hand-made stretching apparatus based on the force gauge and slider system explained in the text. In the former method, we use small samples of width 10 mm and of height $L=18$ mm. In the latter method, we use larger samples of width 150 mm and of height 150 mm. The sample height $L$ determines the strain rate, which is in the range 0.025 to 0.2 s-1 in the velocity jump experiment. In each of plot in Fig. 4 (a) to (d), there are two curves of the same color, which correspond to the data obtained under the same condition but with different sample and thus demonstrate the sample dependence of the results.
In Fig. 4 (a), we show results obtained with the rheometer, where the strain rate (= 0.002 s-1) is rather small compared with those in the jump experiment. The results suggest (1) good reproducibility, (2) small dependence on the stretching direction, and (3) reasonable linearity up to the critical strain $\varepsilon_{c}\sim 0.1$ although strictly linear region is limited in the region $\varepsilon<0.02$. The stress and strain at the breakage tend to be slightly small in the MD direction, which is barely consistent with the fact that molecules tend to be oriented in the TD direction. It is possible that the breakage stress could be clearly larger than that in the TD direction, if the experiment were performed under a better control of the sharpness of the free edge (see the next paragraph).
In Fig. 4 (b), we compare the results obtained by the two methods. The agreement of the results is good in the region $\varepsilon<\varepsilon_{c}$, but the strain and stress at the breakage is significantly larger for larger samples. This difference may be due to the quality of cutting the sample. The free edge of large samples is cut in the factory at the fabrication in the company, while that of smaller samples is cut in our laboratory. The sharpness of the edge is lower for smaller samples used for the rheometer, which leads to smaller strains at the breakage.
In Fig. 4 (c) and (d), we show results obtained with the hand-made setup, which show the dependence of the curve on the stretching speed, where the strain rate is in the range 0.002 to 0.07 s-1. Although the strain rate tends to be small compared with that in the jump experiment, we see the sample tends to become hard as the rate increases, which is consistent with rheological measurements (see below).
Rheological properties — In Fig. 5, we show results of rheological measurements performed by the rheometer (MCR 302, Anton Paar) at the frequency $f=1$ and 30 Hz. Although PVDC is a crystalline polymer from the monomer structure, the degree of crystallization of the sample may be relatively low and thus transparent. As a result, we see the first relaxation of amorphous structure and the second of crystalline structure in the plot, which respectively correspond to the first and second peak of $\tan\delta$.  Judging from the first peak of $\tan\delta$, the glass transition occurs at $T\simeq$ $13$ C and $f=1$ Hz and at $T\simeq$ $22$ C and $f=30$ Hz. We conclude that the characteristic time scale for the glass transition is $\tau\simeq 1/f\simeq$ tens of milliseconds at the room temperature ($T\simeq$ $25$ C). The wavy feature observed in the plots of $E^{\prime\prime}$ and $\tan\delta$ in the region T=50T=-50 to 15-15 C at $f=1$ Hz is due to insufficient torque and the sharp drop of $^{\prime}, E′′E^{\prime\prime}$, and $\tan\delta$ is due to inertia. These technical artifacts could be ameliorated but, practically, the conclusion on the characteristic time-scale of the glass transition would not be changed.